The length of a rectangle is 2ft longer than its width. Find the dimensions of the rectangle such that the perimeter of the rectangle is 845ft
Answer by jim_thompson5910(35256)

P = 2L + 2W

845 = 2(W+2) + 2W

845 = 2W+4 + 2W

845 = 4W+4

845-4 = 4W

841 = 4W

841/4 = W

W = 841/4

Now that you know W, use it to find L

Remember that L = W+2